# REGULAR ARTICLE

### v1

@@ 1 Introduction
— Albert Einstein [1]
The control of fire and the mastery of the wheel have been regarded as two of humanity’s greatest inventions. Even today, it is difficult to imagine building any complex structure without the wheel or the underlying principle of circular motion on an axis. However, it was not until 1763 that, while repairing Newcomen engines, James Watt combined the concepts of fire and the wheel to create the steam engine [2]. This innovation harnessed the energy of coal flames to produce rotary motion, powering locomotives, factory machinery and setting the stage for the Industrial Revolution. To pursuit the desire to optimise these widely used heat engines, a remarkable group of people such as Sadi Carnot, James Joule, Rudolf Clausius, William Thompson, James Clerk Maxwell and Willard Gibbs laid the foundations of thermodynamics [3]. Though initially formulated for equilibrium and reversible processes, thermodynamics eventually came to incorporate the crucial roles of fluctuations and irreversibility — a topic of the present article. From the 1900s onwards, we further unveiled the atomic and molecular world, giving rise to quantum mechanics. More recently, our curiosity has turned to the question: ‘Are quantum thermodynamic machines better than their classical counterparts?’ [4] Just as the combination of fire and the wheel marked the beginning of the technological revolution, the harnessing of quantum resources may herald a new era. This review aims to briefly trace that remarkable journey.
@@
2 Fluctuations in Equilibrium and Near-equilibrium Thermodynamics
@@
— Arthur Eddington [5]
From the middle of the twentieth century, our understanding of thermodynamics has changed dramatically. While classical thermodynamics relies on equilibrium properties, increasing importance is now given towards the study of fluctuations, instabilities, and evolutionary processes at all levels, from chemistry, biology to cosmology [6]. Almost all processes in nature are irreversible, in which time-reversal symmetry is broken. The distinction between reversible and irreversible processes was introduced into thermodynamics through the concept of “entropy” — the arrow of time, as Arthur Eddington called it [5]. Unlike Newtonian mechanics, which originated in the study of celestial bodies, equilibrium thermodynamics — as emphasised in the introduction — arose during the Industrial Revolution as a phenomenological theory of macroscopic objects with a much more practical purpose: replacing manual labour with efficient machines by extracting useful work from the motive power of fire [7]. To understand the thermodynamic behaviour of a system, it is therefore imperative to examine how it exchanges energy with its surroundings in the form of heat and work. Thus, the foundation of classical thermodynamics rests on two main pillars: energy and entropy [8]. Accordingly, thermodynamic systems are classified into three categories: isolated, closed, and open. In doing so, thermodynamics introduces the notion of temperature — a concept which was absent from classical mechanics. If the temperature of a system is not uniform everywhere, heat will flow until the entire system reaches a single temperature. Such a macroscopic state, characterized by the absence of change on the average, is called an equilibrium state of the system. In particular, traditional thermodynamics focuses on average properties of macroscopic systems and disregards fluctuations. Since thermodynamic systems are typically large, thermal fluctuations are negligible; however, as the system size decreases, both thermal as well as quantum fluctuations become inevitable and must be taken into account [9]. So, the description of non-equilibrium process, requires not only the averages but also the deviations of various thermodynamic quantities from its equilibrium values.
In its original form, an irreversible transformation was characterized by Clausius [10] in terms of
@@
where $S$ is the entropy of the final state and $S_{0}$ is the entropy of the initial state. According to him, “The magnitude $N$ thus determines the uncompensated transformation (uncompensirte Verwandlung)” [11]. It represents the entropy produced by irreversible processes, which, by virtue of the second law of thermodynamics, can only be positive [12]. This production of entropy within the system establishes an arrow of time, distinguishing the future from the past. This stands in sharp contrast to the laws of mechanics — both classical and quantum — which are time symmetric and possess no intrinsic time-irreversibility in their fundamental equations. Thus, processes ruled out by the second law of thermodynamics, such as the spontaneous flow of gas molecules from a region of lower concentration to one of higher concentration, don’t contradict with the laws of mechanics [6]. This leads to a fundamental question: How can irreversible macroscopic processes, such as the flow of heat from higher to lower temperatures, emerge from the reversible microscopic dynamics of atoms and molecules? The answer lies in the foundations of statistical mechanics, which connect the probabilistic description of microscopic states with the macroscopic behaviour of thermodynamic systems [13].
To address these pertinent questions, Ludwig Boltzmann was first to introduce an extraordinary relation — engraved on his memorial in the Central Cemetery in Vienna — that entropy is a logarithmic measure of the number of microstates: [14]
@@
Expanding $S$ in a Taylor series around the equilibrium value $A_{i}^{0}$ and retaining terms up to second order, we obtain [6]
@@
Later, Albert Einstein prescribed a formula by rewriting Boltzmann’s idea in reverse: [15]
@@
$P(A_{i})=\frac{W}{W_{0}}=e^{\frac{\Delta S}{k_{B}}}=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\bigg[-\frac{(A_{i}-A_{i}^{0})^{2}}{2\sigma^{2}}\bigg],$ (8)
@@
and is positive definite [6]. As an example, if we calculate the fluctuation in volume for an ideal gas [16], the relative fluctuation tends to zero for large $N$:
@@
Now, for multiple variables, we introduce the Onsager coordinates $\alpha_{i}=A_{i}-A_{i}^{0}$ (i=1,2,)(i=1,2,...), which represent the fluctuations of different thermodynamic variables around their equilibrium values. Proceeding as before, we expand the entropy, $S=f(\{A_{i}\})$, up to linear-order terms in $\alpha_{i}$s’: [6, 16]
@@
Since the system is in equilibrium, this change is negative, as before. So, the equilibrium state is stable against any perturbation that decreases entropy. Conversely, if fluctuation grows, the system is not in equilibrium. This principle is known as Gibbs stability theory [6].
In response to a fluctuation that decreases entropy from its maximum equilibrium value, there will be irreversible processes producing entropy ($\Delta_{i}S$) that spontaneously drive the system back to equilibrium [6]. At equilibrium, these processes vanish and 19th-century thermodynamics was largely focused on such idealised reversible transformations. Recognising the connection between entropy and irreversibility, Pierre Duhem began developing a formalism that was later completed by Prigogine [12] and Onsager [17, 18]. In the modern approach, entropy change is calculated in terms of variables that characterise non-equilibrium processes. Even if a system is not in global equilibrium, thermodynamic quantities such as temperature, concentration, pressure, and internal energy remain well-defined locally, i.e., intensive parameters (temperature, pressure, etc.) are well defined within small volume elements, while extensive variables (entropy, internal energy, etc.) are expressed in terms of their corresponding densities. This is the assumption of local equilibrium, which holds pretty well for most physical and chemical systems. [12, 6]
Modern formalism due to Prigogine, begins with the infinitesimal change in entropy, written as [12, 6]
@@
where $d_{e}S$ is the entropy change due to the exchange of matter and energy with the environment, and $d_{i}S$ is the entropy change due to the uncompensated transformation, i.e., the entropy produced by irreversible transformations. This in turn can be described by the sum of individual thermodynamic forces (such as gradients of temperature, concentration, etc.) and their corresponding thermodynamic fluxes (such as heat flow, diffusion, etc.). The rate of entropy production can thus be expressed as [8, 6]
@@
where $F_{k}$ and $J_{k}$ are the respective thermodynamic forces and fluxes. The entropy production rate and the entropy production itself are always positive ($\dot{\Sigma}>0,\;d_{i}S>0$). This statement is stronger and more general than the classical formulation that the entropy of an isolated system can only increase. This insight by Prigogine earned him the Nobel Prize in Chemistry in 1977 [19].
To obtain Eq. (14) for the $\dot{\Sigma}$ associated with fluctuations in system variables, the thermodynamic force can be written as, $F_{k}=\frac{\partial\Delta_{i}S}{\partial\alpha_{k}}$, conjugate to the thermodynamic flux $J_{k}=\frac{\partial\alpha_{k}}{\partial t}$. If the deviations of the forces from their equilibrium values are small, the fluxes can be expressed as linear functions of the forces, $J_{k}=\sum_{j}L_{kj}F_{j}$, where $L_{kj}$ are constants known as kinetic coefficients. Thus, the rate of entropy production can be calculated as [6]
@@
Here, $\mathbf{L}$ is a positive definite matrix whose elements satisfy the symmetry condition $L_{kj}=L_{jk}$. These reciprocal relations are linked to thermodynamic cross-effects and were first observed by Lord Kelvin and others in the study of thermoelectric phenomena (such as the Seebeck (1821) and Peltier (1834) effects in metal wires) [6, 20]. At that time, they were regarded merely as conjectures. A rigorous mathematical derivation was later provided by Lars Onsager in 1931, after which they became known as the Onsager reciprocal relations [17, 18]. The foundation of Onsager’s theory rests on the properties of steady-state fluctuations:
@@
The first equality reflects time-reversal symmetry, while the second expresses time-translation invariance. Taken together, these properties — often referred to as microscopic reversibility or the principle of detailed balance — lead directly to the reciprocal relations $L_{kj}=L_{jk}$. The importance of this result is so profound that it is regarded as “fourth law of thermodynamics,” [21] a contribution for which Onsager was awarded the Nobel Prize in Chemistry in 1968 [22].
We conclude this section by highlighting the deep connection between entropy and fluctuations in equilibrium and near-equilibrium regimes. This connection explains why equilibrium thermodynamic properties are primarily governed by average values and how microscopic fluctuation symmetries give rise to thermodynamic cross-effects at the macroscopic scale. To fully comprehend how irreversibility emerges at macroscopic length scales from time-reversible microscopic dynamics, one must turn to the field of stochastic thermodynamics [13], which came to the forefront in the 1990s.
3 Emergence of Stochastic Thermodynamics and Fluctuation Theorems
@@
— Christopher Jarzynski [23]
Over the past three decades, stochastic thermodynamics has become a leading framework for understanding diverse non-equilibrium phenomena, ranging from molecular motors and cellular reactions to colloids and nanoscale systems [24, 25, 26, 27]. Often, thermodynamics on this scale refers as nanoscale thermodynamics [28], and researchers are now attempting to extend this concept to the quantum realm, which will be discussed in the following section [29]. Broadly speaking, stochastic thermodynamics describes mesoscopic nonequilibrium systems in contact with equilibrium heat bath. Here, the system interacts randomly with the reservoir, so that at time $t$ it may be found in a state $x$ with probability $P(x;t)$. In such small systems, fluctuations play a pivotal role in controlling the dynamics of the system, and thermodynamic quantities such as heat, work, and entropy themselves fluctuate. Unlike in the classical case, these quantities don’t take on single fixed values but are instead described by probability distributions, with mean values serving as the relevant averages. Remarkably, these distributions for thermodynamic observables fulfil some symmetry relations known as fluctuation theorems (FTs) [27]. Here, we will discuss a few such relations.
3.0.1 Evans-Searles fluctuation theorem (1994):
Evans and Searles proposed the first fluctuation relation related to general dissipation in simulations of sheared fluids [30]. Here, we present a general version of the theorem within Hamiltonian dynamics. Consider an isolated Hamiltonian system of interacting particles. The state of the system is represented by a phase-space vector containing the coordinates and canonical momenta of all particles, in an Avogadro-dimensional space, denoted as $\Gamma\equiv\{\textbf{q}_{i},\textbf{p}_{i}\}$. The Hamilton’s equations for such an autonomous system are given by
@@ 3.0.1 Evans-Searles fluctuation theorem (1994):
$\Omega_{\tau}(\Gamma_{0})=\ln\bigg(\frac{P(d\Gamma_{0},0)}{P(d\Gamma_{0}^{*},0)}\bigg)=\ln\bigg(\frac{f(d\Gamma_{0},0)}{f(d\Gamma_{0}^{*},0)}\bigg)-\int^{\tau}_{0}\Lambda dt,$ (27)
which quantifies the deviation from equilibrium. For $\Omega_{\tau}(\Gamma_{0})$ to be well defined over the phase space domain $D$, the system must be ergodically consistent over $D$ [13]. It is evident that $\Omega_{\tau}(\Gamma_{0})=0$ at equilibrium. Since $\Omega_{\tau}(\Gamma_{0})$ characterizes the reversibility of a set of trajectories, we can evaluate the relative probability of the dissipation function taking opposite values:
@@
where we used the key property of the odd parity of the dissipation function, i.e., $\Omega_{\tau}(\Gamma_{0})=-\Omega_{\tau}(\Gamma_{0}^{*})$, which follows directly from its definition in Eq. (27). Equation (28) is known as the Evans–Searles fluctuation theorem (ESFT) [30]. This implies that dissipation functions taking positive values for forward irreversible trajectories are more probable than for the corresponding reverse antitrajectories. Thus, the dissipation function can be viewed as a measure of the temporal asymmetry inherent in bundles of trajectories originating from an initial distribution of states. In defining the dissipation function in Eq. (27), which characterizes deviations from equilibrium, we did not restrict ourselves to the near-equilibrium situation. As a consequence, the ESFT derived here applies to nonequilibrium fluctuations arbitrarily far from equilibrium and to systems of any size, without requiring the classical thermodynamic limit. Most importantly, the ESFT demonstrates that even deterministic, time-reversible equations of motion can give rise to irreversibility [13]. The ESFT has also been verified experimentally, for instance by Wang et al. (2002) [31] and Reid et al. (2004) [32].
One can further evaluate the phase-space average of the dissipation function as $\langle\exp(-\Omega_{\tau})\rangle=1$. Since $\exp(-\Omega_{\tau})$ is a convex function, Jensen’s inequality gives [27]
@@
which implies $\langle\Omega_{\tau}\rangle\geq 0$. This shows that although the dissipation function defined in Eq. (27) can take positive, negative, or zero values, its average is always nonnegative. This is the statement of the second law of thermodynamics, highlighting the link between microscopic reversible dynamics and macroscopic irreversible behavior, and providing a resolution of the Loschmidt paradox [13]. In the following, we extend the FTs to more practical thermodynamic quantities.
3.0.2 Jarzynski equality (1997):
Jarzynski developed a more practical relation that allows us to relate the work $W$ and the free energy change $\Delta F$ for a non-quasistatic process between two sates of equilibrium. It is an equality and hence contains more information than classical thermodynamics, which only connects them through an inequality $W\geq\Delta F$ [33, 34].

### v2

# Fluctuations and Irreversibility: Historical and Modern Perspectives

@@ I Introduction
— Albert Einstein einstein1926
The control of fire and the mastery of the wheel have been regarded as two of humanity’s greatest inventions. Even today, it is difficult to imagine building any complex structure without the wheel or the underlying principle of circular motion on an axis. However, it was not until 1763 that, while repairing Newcomen engines, James Watt combined the concepts of fire and the wheel to create the steam engine rao2011history . This innovation harnessed the energy of coal flames to produce rotary motion, powering locomotives, factory machinery and setting the stage for the Industrial Revolution. To pursuit the desire to optimise these widely used heat engines, a remarkable group of people such as Sadi Carnot, James Joule, Rudolf Clausius, William Thompson, James Clerk Maxwell and Willard Gibbs laid the foundations of thermodynamics muller2007history . Though initially formulated for equilibrium and reversible processes, thermodynamics eventually came to incorporate the crucial roles of fluctuations and irreversibility — a topic of the present article. From the 1900s onwards, we further unveiled the atomic and molecular world, giving rise to quantum mechanics. More recently, our curiosity has turned to the question: ‘Are quantum thermodynamic machines better than their classical counterparts?’ ghosh2019are Just as the combination of fire and the wheel marked the beginning of the technological revolution, the harnessing of quantum resources may herald a new era. This review aims to briefly trace that remarkable journey.
@@
II Fluctuations in Equilibrium and Near-equilibrium Thermodynamics
@@
— Arthur Eddington eddington1928book
From the middle of the twentieth century, our understanding of thermodynamics has changed dramatically. While classical thermodynamics relies on equilibrium properties, increasing importance is now given towards the study of fluctuations, instabilities, and evolutionary processes at all levels, from chemistry, biology to cosmology kondepudi2015book . Almost all processes in nature are irreversible, in which time-reversal symmetry is broken. The distinction between reversible and irreversible processes was introduced into thermodynamics through the concept of “entropy” — the arrow of time, as Arthur Eddington called it eddington1928book . Unlike Newtonian mechanics, which originated in the study of celestial bodies, equilibrium thermodynamics — as emphasised in the introduction — arose during the Industrial Revolution as a phenomenological theory of macroscopic objects with a much more practical purpose: replacing manual labour with efficient machines by extracting useful work from the motive power of fire carnot1824book . To understand the thermodynamic behaviour of a system, it is therefore imperative to examine how it exchanges energy with its surroundings in the form of heat and work. Thus, the foundation of classical thermodynamics rests on two main pillars: energy and entropy callen1985book . Accordingly, thermodynamic systems are classified into three categories: isolated, closed, and open. In doing so, thermodynamics introduces the notion of temperature — a concept which was absent from classical mechanics. If the temperature of a system is not uniform everywhere, heat will flow until the entire system reaches a single temperature. Such a macroscopic state, characterized by the absence of change on the average, is called an equilibrium state of the system. In particular, traditional thermodynamics focuses on average properties of macroscopic systems and disregards fluctuations. Since thermodynamic systems are typically large, thermal fluctuations are negligible; however, as the system size decreases, both thermal as well as quantum fluctuations become inevitable and must be taken into account binder2019thermodynamics . So, the description of non-equilibrium process, requires not only the averages but also the deviations of various thermodynamic quantities from its equilibrium values.
In its original form, an irreversible transformation was characterized by Clausius clausius1891book in terms of
@@
where $S$ is the entropy of the final state and $S_{0}$ is the entropy of the initial state. According to him, “The magnitude $N$ thus determines the uncompensated transformation (uncompensirte Verwandlung)” clausius1879book . It represents the entropy produced by irreversible processes, which, by virtue of the second law of thermodynamics, can only be positive prigogine1961book . This production of entropy within the system establishes an arrow of time, distinguishing the future from the past. This stands in sharp contrast to the laws of mechanics — both classical and quantum — which are time symmetric and possess no intrinsic time-irreversibility in their fundamental equations. Thus, processes ruled out by the second law of thermodynamics, such as the spontaneous flow of gas molecules from a region of lower concentration to one of higher concentration, don’t contradict with the laws of mechanics kondepudi2015book . This leads to a fundamental question: How can irreversible macroscopic processes, such as the flow of heat from higher to lower temperatures, emerge from the reversible microscopic dynamics of atoms and molecules? The answer lies in the foundations of statistical mechanics, which connect the probabilistic description of microscopic states with the macroscopic behaviour of thermodynamic systems evans2016fundamentals .
To address these pertinent questions, Ludwig Boltzmann was first to introduce an extraordinary relation — engraved on his memorial in the Central Cemetery in Vienna — that entropy is a logarithmic measure of the number of microstates: lebowitz1993boltzmann
@@
Expanding $S$ in a Taylor series around the equilibrium value $A_{i}^{0}$ and retaining terms up to second order, we obtain kondepudi2015book
@@
Later, Albert Einstein prescribed a formula by rewriting Boltzmann’s idea in reverse: sommerfeld1964book
@@
$P(A_{i})=\frac{W}{W_{0}}=e^{\frac{\Delta S}{k_{B}}}=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp[-\frac{(A_{i}-A_{i}^{0})^{2}}{2\sigma^{2}}\bigg],$ (8)
@@
and is positive definite kondepudi2015book . As an example, if we calculate the fluctuation in volume for an ideal gas dsray2015note , the relative fluctuation tends to zero for large $N$:
@@
Now, for multiple variables, we introduce the Onsager coordinates $\alpha_{i}=A_{i}-A_{i}^{0}$ (i=1,2,)(i=1,2,...), which represent the fluctuations of different thermodynamic variables around their equilibrium values. Proceeding as before, we expand the entropy, $S=f(\{A_{i}\})$, up to linear-order terms in $\alpha_{i}$s’: kondepudi2015book ; dsray2015note
@@
Since the system is in equilibrium, this change is negative, as before. So, the equilibrium state is stable against any perturbation that decreases entropy. Conversely, if fluctuation grows, the system is not in equilibrium. This principle is known as Gibbs stability theory kondepudi2015book .
In response to a fluctuation that decreases entropy from its maximum equilibrium value, there will be irreversible processes producing entropy ($\Delta_{i}S$) that spontaneously drive the system back to equilibrium kondepudi2015book . At equilibrium, these processes vanish and 19th-century thermodynamics was largely focused on such idealised reversible transformations. Recognising the connection between entropy and irreversibility, Pierre Duhem began developing a formalism that was later completed by Prigogine prigogine1961book and Onsager onsager1931reciprocal-I ; onsager1931reciprocal-II . In the modern approach, entropy change is calculated in terms of variables that characterise non-equilibrium processes. Even if a system is not in global equilibrium, thermodynamic quantities such as temperature, concentration, pressure, and internal energy remain well-defined locally, i.e., intensive parameters (temperature, pressure, etc.) are well defined within small volume elements, while extensive variables (entropy, internal energy, etc.) are expressed in terms of their corresponding densities. This is the assumption of local equilibrium, which holds pretty well for most physical and chemical systems. prigogine1961book ; kondepudi2015book
Modern formalism due to Prigogine, begins with the infinitesimal change in entropy, written as prigogine1961book ; kondepudi2015book
@@
where $d_{e}S$ is the entropy change due to the exchange of matter and energy with the environment, and $d_{i}S$ is the entropy change due to the uncompensated transformation, i.e., the entropy produced by irreversible transformations. This in turn can be described by the sum of individual thermodynamic forces (such as gradients of temperature, concentration, etc.) and their corresponding thermodynamic fluxes (such as heat flow, diffusion, etc.). The rate of entropy production can thus be expressed as callen1985book ; kondepudi2015book
@@
where $F_{k}$ and $J_{k}$ are the respective thermodynamic forces and fluxes. The entropy production rate and the entropy production itself are always positive ($\dot{\Sigma}>0,\;d_{i}S>0$). This statement is stronger and more general than the classical formulation that the entropy of an isolated system can only increase. This insight by Prigogine earned him the Nobel Prize in Chemistry in 1977 prigogine1977time .
To obtain Eq. (14) for the $\dot{\Sigma}$ associated with fluctuations in system variables, the thermodynamic force can be written as, $F_{k}=\frac{\partial\Delta_{i}S}{\partial\alpha_{k}}$, conjugate to the thermodynamic flux $J_{k}=\frac{\partial\alpha_{k}}{\partial t}$. If the deviations of the forces from their equilibrium values are small, the fluxes can be expressed as linear functions of the forces, $J_{k}=\sum_{j}L_{kj}F_{j}$, where $L_{kj}$ are constants known as kinetic coefficients. Thus, the rate of entropy production can be calculated as kondepudi2015book
@@
Here, $\mathbf{L}$ is a positive definite matrix whose elements satisfy the symmetry condition $L_{kj}=L_{jk}$. These reciprocal relations are linked to thermodynamic cross-effects and were first observed by Lord Kelvin and others in the study of thermoelectric phenomena (such as the Seebeck (1821) and Peltier (1834) effects in metal wires) kondepudi2015book ; gupt2024graph . At that time, they were regarded merely as conjectures. A rigorous mathematical derivation was later provided by Lars Onsager in 1931, after which they became known as the Onsager reciprocal relations onsager1931reciprocal-I ; onsager1931reciprocal-II . The foundation of Onsager’s theory rests on the properties of steady-state fluctuations:
@@
The first equality reflects time-reversal symmetry, while the second expresses time-translation invariance. Taken together, these properties — often referred to as microscopic reversibility or the principle of detailed balance — lead directly to the reciprocal relations $L_{kj}=L_{jk}$. The importance of this result is so profound that it is regarded as “fourth law of thermodynamics,” deffner2019quantumthermo a contribution for which Onsager was awarded the Nobel Prize in Chemistry in 1968 onsager1968the .
We conclude this section by highlighting the deep connection between entropy and fluctuations in equilibrium and near-equilibrium regimes. This connection explains why equilibrium thermodynamic properties are primarily governed by average values and how microscopic fluctuation symmetries give rise to thermodynamic cross-effects at the macroscopic scale. To fully comprehend how irreversibility emerges at macroscopic length scales from time-reversible microscopic dynamics, one must turn to the field of stochastic thermodynamics evans2016fundamentals , which came to the forefront in the 1990s.
III Emergence of Stochastic Thermodynamics and Fluctuation Theorems
@@
— Christopher Jarzynski jarzynski2015diverse
Over the past three decades, stochastic thermodynamics has become a leading framework for understanding diverse non-equilibrium phenomena, ranging from molecular motors and cellular reactions to colloids and nanoscale systems udo2012stochastic ; peliti2021stochastic ; limmer2024statistical ; seifert2025stochastic . Often, thermodynamics on this scale refers as nanoscale thermodynamics broedersz2022twenty , and researchers are now attempting to extend this concept to the quantum realm, which will be discussed in the following section benenti2017fundamental . Broadly speaking, stochastic thermodynamics describes mesoscopic nonequilibrium systems in contact with equilibrium heat bath. Here, the system interacts randomly with the reservoir, so that at time $t$ it may be found in a state $x$ with probability $P(x;t)$. In such small systems, fluctuations play a pivotal role in controlling the dynamics of the system, and thermodynamic quantities such as heat, work, and entropy themselves fluctuate. Unlike in the classical case, these quantities don’t take on single fixed values but are instead described by probability distributions, with mean values serving as the relevant averages. Remarkably, these distributions for thermodynamic observables fulfil some symmetry relations known as fluctuation theorems (FTs) seifert2025stochastic . Here, we will discuss a few such relations.
III.0.1 Evans-Searles fluctuation theorem (1994):
Evans and Searles proposed the first fluctuation relation related to general dissipation in simulations of sheared fluids evans1994equilibrium . Here, we present a general version of the theorem within Hamiltonian dynamics. Consider an isolated Hamiltonian system of interacting particles. The state of the system is represented by a phase-space vector containing the coordinates and canonical momenta of all particles, in an Avogadro-dimensional space, denoted as $\Gamma\equiv\{\textbf{q}_{i},\textbf{p}_{i}\}$. The Hamilton’s equations for such an autonomous system are given by
@@ III.0.1 Evans-Searles fluctuation theorem (1994):
$\Omega_{\tau}(\Gamma_{0})=\ln\bigg(\frac{P(d\Gamma_{0},0)}{P(d\Gamma_{0}^{*},0)}\bigg.)=\ln\bigg(\frac{f(d\Gamma_{0},0)}{f(d\Gamma_{0}^{*},0)}\bigg.)-\int^{\tau}_{0}\Lambda dt,$ (27)
which quantifies the deviation from equilibrium. For $\Omega_{\tau}(\Gamma_{0})$ to be well defined over the phase space domain $D$, the system must be ergodically consistent over $D$ evans2016fundamentals . It is evident that $\Omega_{\tau}(\Gamma_{0})=0$ at equilibrium. Since $\Omega_{\tau}(\Gamma_{0})$ characterizes the reversibility of a set of trajectories, we can evaluate the relative probability of the dissipation function taking opposite values:
@@
where we used the key property of the odd parity of the dissipation function, i.e., $\Omega_{\tau}(\Gamma_{0})=-\Omega_{\tau}(\Gamma_{0}^{*})$, which follows directly from its definition in Eq. (27). Equation (28) is known as the Evans–Searles fluctuation theorem (ESFT) evans1994equilibrium . This implies that dissipation functions taking positive values for forward irreversible trajectories are more probable than for the corresponding reverse antitrajectories. Thus, the dissipation function can be viewed as a measure of the temporal asymmetry inherent in bundles of trajectories originating from an initial distribution of states. In defining the dissipation function in Eq. (27), which characterizes deviations from equilibrium, we did not restrict ourselves to the near-equilibrium situation. As a consequence, the ESFT derived here applies to nonequilibrium fluctuations arbitrarily far from equilibrium and to systems of any size, without requiring the classical thermodynamic limit. Most importantly, the ESFT demonstrates that even deterministic, time-reversible equations of motion can give rise to irreversibility evans2016fundamentals . The ESFT has also been verified experimentally, for instance by Wang et al. (2002) wang2002experimental and Reid et al. (2004) reid2004reversibility .
One can further evaluate the phase-space average of the dissipation function as $\langle\exp(-\Omega_{\tau})\rangle=1$. Since $\exp(-\Omega_{\tau})$ is a convex function, Jensen’s inequality gives seifert2025stochastic
@@
which implies $\langle\Omega_{\tau}\rangle\geq 0$. This shows that although the dissipation function defined in Eq. (27) can take positive, negative, or zero values, its average is always nonnegative. This is the statement of the second law of thermodynamics, highlighting the link between microscopic reversible dynamics and macroscopic irreversible behavior, and providing a resolution of the Loschmidt paradox evans2016fundamentals . In the following, we extend the FTs to more practical thermodynamic quantities.
III.0.2 Jarzynski equality (1997):
Jarzynski developed a more practical relation that allows us to relate the work $W$ and the free energy change $\Delta F$ for a non-quasistatic process between two sates of equilibrium. It is an equality and hence contains more information than classical thermodynamics, which only connects them through an inequality $W\geq\Delta F$ jarzynski1997nonequilibrium ; jarzynski2010equalities .